# Explainable Linear and Generalized Linear Models by the Predictions Plot

### v2

@@ Abstract
Many statistics courses cover multiple linear regression, and present students with the formula of a prediction using the input variables, slopes, and an intercept. But is it really easy to see which terms have the largest effect, or to explain why the prediction of a specific case is unusually high or low? To assist with this the so-called predictions plot is proposed. Its simplicity makes it easy to interpret, and it combines much information. Its main benefit is that it helps explainability of the prediction formula as it is, without depending on how the formula was derived. The input variables can be numerical or categorical; interaction terms are also handled, and the model can be linear or generalized linear. Another display is proposed to visualize correlations between prediction terms, in a way that is tailored for this setting.
@@ 2 Spread and orientation of prediction terms
To illustrate this we analyze the well-known Titanic dataset which contains information on the 1309 passengers of the RMS Titanic. For its history and visualizations see Friendly et al. (2019). The data are freely available from https://www.kaggle.com/c/titanic/data and the R package classmap (Raymaekers and Rousseeuw, 2023). The binary response variable indicates whether the passenger survived or was a casualty. We want to predict survival by a logistic regression on the variables pclass, sex, age, sibsp, and parch. Here pclass is the cabin class, sex is M or F, and the passenger’s age is in years. The variables sibsp and parch count the number of siblings+spouses and parents+children aboard.
@@
It is hoped that the way the predictions plot visually combines much information provides helpful insight in the output of a regression. Figures 2 and 3 were obtained by the function predsplot in the Supplementary Material using the single R command predsplot(fit) where fit is an output object of the function lm() or glm(). The function predsplot figures out the rest. Of course there are many options: the maximal number of prediction terms to be shown, graphical parameters for colors and line widths, and so on. There is also an argument to specify whether to display histograms or estimated densities. For densities there are arguments that input the desired bandwidth as well. Here we used the defaults of stats::density(). The formula of the regression may include logarithms of variables and interaction terms, such as $y\sim x_{1}+\log(x_{2})+x_{3}:x_{4}$ as illustrated in Section A of the Supplementary Material.
@@ 4 Correlations between prediction terms
The off-diagonal part of the display consists of rectangles, which also have a meaningful interpretation. The area of such a rectangle is the product of the standard deviations of its prediction terms, and when we multiply that with the value of the correlation (represented by its color) we obtain the covariance between the prediction terms. So a dark color is more important in a larger rectangle (e.g. formed by months and amount) than in a smaller rectangle (e.g. formed by rate and amount). The proposed display thus depicts the covariance matrix between the prediction terms, which is relevant in a regression. Such a display is not recommended for a covariance matrix between variables in different units, but here all prediction terms are in the same units (those of the total linear prediction).
The R function predscor which produces this display has various options, such as sorting the prediction terms by decreasing standard deviation, plotting the absolute values of the correlations, or making the area of the diagonal cells proportional to the standard deviation of the prediction terms rather than their variance.
@@
Figure 7 shows the resulting predictions plot. It looks rather suspicious, because the artificial variables x1 and x2 are more spread out than the total prediction on the right, which is the sum of these and other terms. Moreover, x1 points downward and x2 points upward! They partly cancel each other. The correlation display in Figure 8 also looks unusual, with the big dark blue rectangles between x1 and x2. The correlation between x1 and x2 that was close to 1 has flipped sign between their prediction terms, again reflecting the canceling effect. In general, prediction terms are easiest to interpret when they only have small correlations with each other.
@@ 5 Related methods
A coefficient plot, described for instance by Arel-Bundock (2022), does look at several input variables at the same time. But it does not plot the distributions of the input variables or the prediction. Instead it plots the regression coefficients associated with the input variables, that need to be numerical, as well as their confidence intervals. The coefficients are plotted as points, and the confidence intervals as line segments. The coefficients can be made comparable to each other if the input variables are first standardized, say to zero mean and unit standard deviation. The coefficient plot can summarize an entire model, but in a way very different from a predictions plot.
@@ 6 Discussion
Software Availability. The R code with an example script that reproduces the figures is available at https://wis.kuleuven.be/statdatascience/code/predsplot_r_code.zip . It will be incorporated in an R package on CRAN.
@@ Appendix A More figures of the Top Gear data
It is a simple example because all of the input variables are numeric. The question was, which of these has the biggest effect on the prediction. Using predsplot() we can easily answer this:
> gp = predsplot(fit, main = "Top Gear data")
> print(gp\$predsummary, row.names = F)
@@
We see that the prediction term from disp has the largest standard deviation, followed by topspeed, whereas length only has a tiny effect. The resulting predictions plot visualizes this:
> plot(gp\$p)
We clearly see that the range of the prediction term of length is tiny. This is confirmed by the new display of the correlations and standard deviations of the prediction terms:
@@
> gp = predsplot(fit, main = main)
> plot(gp\$p)
yielding Figure 2, and we get the standard deviations of the prediction terms by
> print(gp\$predsummary, row.names = F)
@@
> plot(predsplot(fit, main = main, casetoshow = car, displaytype = "density",
                 xlab = paste0("prediction for ", car))\$p)
@@
> plot(predsplot(fit, main = main, casetoshow = car, staircase = TRUE,
               xlab = paste0("prediction for ", car))\$p)
@@
To illustrate that the code can handle logarithms and interactions in the formula we run
@@ Appendix B More figures of the Titanic data
> plot(predsplot(fit, main = main, casetoshow=1)\$p)
> plot(predsplot(fit, main = main, casetoshow=2, staircase = TRUE)\$p)
@@ Appendix C More figures of the German credit data
> gp = predsplot(fit, main = main, casetoshow = newc, staircase = TRUE)
> plot(gp\$p)
@@
In the output we see that the new predicted probability is about $89\%$ :
> print(gp\$casesummary, row.names = F)
@@
> plot(predsplot(fit, main = main, casetoshow = 1,
                 displaytype = "density", profile = TRUE)\$p)

### v3

@@ Abstract
Multiple linear regression is a basic statistical tool, yielding a prediction formula with the input variables, slopes, and an intercept. But is it really easy to see which terms have the largest effect, or to explain why the prediction of a specific case is unusually high or low? To assist with this the so-called predictions plot is proposed. Its simplicity makes it easy to interpret, and it combines much information. Its main benefit is that it helps explainability of the prediction formula as it is, without depending on how the formula was derived. The input variables can be numerical or categorical. Interaction terms are also handled, and the model can be linear or generalized linear. Another display is proposed to visualize correlations and covariances between prediction terms, in a way that is tailored for this setting.
@@ 2 Spread and orientation of prediction terms
To illustrate this we analyze the well-known Titanic dataset which contains information on the 1309 passengers of the RMS Titanic. For its history and visualizations see Friendly et al. (2019). The data are freely available from https://www.kaggle.com/c/titanic/data and the R package classmap (Raymaekers and Rousseeuw, 2025). The binary response variable indicates whether the passenger survived or was a casualty. We want to predict survival by a logistic regression on the variables pclass, sex, age, sibsp, and parch. Here pclass is the cabin class, sex is M or F, and the passenger’s age is in years. The variables sibsp and parch count the number of siblings+spouses and parents+children aboard.
@@
It is hoped that the way the predictions plot visually combines much information provides helpful insight in the output of a regression. Figures 2 and 3 were obtained by the function predsplot in the R package classmap on CRAN, using the single command predsplot(fit) where fit is an output object of the function lm() or glm(). The function predsplot figures out the rest. Of course there are many options: the maximal number of prediction terms to be shown, graphical parameters for colors and line widths, and so on. There is also an argument to specify whether to display histograms or estimated densities. For densities there are arguments that input the desired bandwidth as well. Here we used the defaults of stats::density(). The formula of the regression may include logarithms of variables and interaction terms, such as $y\sim x_{1}+\log(x_{2})+x_{3}:x_{4}$ as illustrated in Section A of the Supplementary Material.
@@ 4 Correlations between prediction terms
The off-diagonal part of the display consists of rectangles, which also have a meaningful interpretation. The area of such a rectangle is the product of the standard deviations of its prediction terms, and when we multiply that with the value of the correlation (represented by its color) we obtain the covariance between the prediction terms. So a dark color is more important in a larger rectangle (e.g. formed by months and amount) than in a smaller rectangle (e.g. formed by rate and amount). The proposed display thus depicts the covariance matrix between the prediction terms, which is relevant in a regression. Such a display would not be meaningful for a covariance matrix between variables in different units, but here all prediction terms are in the same units (those of the total linear prediction).
The argument of the R function predscor() making this display is again a fit produced by lm() or glm(). There are various options, such as sorting the prediction terms by decreasing standard deviation, plotting the absolute values of the correlations, or making the area of the diagonal cells proportional to the standard deviation of the prediction terms rather than their variance.
@@
Figure 7 shows the resulting predictions plot. It looks rather suspicious, because the artificial variables x1 and x2 are more spread out than the total prediction on the right, which is the sum of these and other terms. Moreover, the predictions of x1 and x2 look like mirror images. They partly cancel each other. The correlation display in Figure 8 also looks unusual, with the big dark blue rectangles between x1 and x2. The correlation between x1 and x2 that was close to 1 has flipped sign between their prediction terms, again reflecting the canceling effect. In general, prediction terms are easiest to interpret when they only have small correlations with each other.
@@ 5 Related methods
A coefficient plot, described for instance by Arel-Bundock (2022), does look at several input variables at the same time. But it does not plot the distributions of the input variables or the prediction. Instead it plots the regression coefficients associated with the input variables, that need to be numerical, as well as their confidence intervals. The coefficients are plotted as points, and the confidence intervals as line segments. The coefficients can be made comparable to each other if the input variables are first standardized, say to zero mean and unit standard deviation. The coefficient plot describes the whole model, but in a way very different from the predictions plot.
@@ 6 Discussion
Software Availability. The functions predsplot() and predscor() have been added to the R package classmap (Raymaekers and Rousseeuw, 2025) on CRAN, together with a vignette that reproduces all the figures in the paper and the supplementary text.
@@ Appendix A More figures of the Top Gear data
It is a simple example because all the input variables are numeric. The question was, which of these has the biggest effect on the prediction. Using predsplot() we can easily answer this:
> library("classmap")
> predsplot(fit, main = "Top Gear data")
@@
We see that the prediction term from displ has the largest standard deviation, followed by topspeed, whereas length only has a tiny effect. The resulting predictions plot visualizes this:
We clearly see that the spread of the prediction term of length is tiny. This is confirmed by the new display of the correlations and standard deviations of the prediction terms:
@@
> predsplot(fit, main = main)
yielding Figure 2, and on the terminal we get the standard deviations of the prediction terms:
@@
> predsplot(fit, main = main, casetoshow = car, displaytype = "density",
                 xlab = paste0("prediction for ", car))
@@
> predsplot(fit, main = main, casetoshow = car, staircase = TRUE,
               xlab = paste0("prediction for ", car))
@@
To illustrate that the code can handle transformations and interactions in the formula we run
@@
The code can also deal with logical and character variables. Let us run
[…]
where alarm (alarm system) is TRUE/FALSE, and navig (satellite navigation) is “y”/“n”.
@@ Appendix B More figures of the Titanic data
> predsplot(fit, main = main, casetoshow=1)
> predsplot(fit, main = main, casetoshow=2, staircase = TRUE)
@@ Appendix C More figures of the German credit data
> predsplot(fit, main = main, casetoshow = newc, staircase = TRUE)
@@
In the output on the terminal we see that the new predicted probability is about $89\%$ :
@@
> predsplot(fit, main = main, casetoshow = 1,
            displaytype = "density", profile = TRUE)